BettyIsAwesome:

I'm sorry. I need more help! (:

BettyIsAwesome:

TogaHimiko:

This question relates to absolute value. |6-4| =2 |3-8|= -5 ur going to turn that negative 5 into a postive 5 then 2-5= -3 so the answer will be -3

TogaHimiko:

@togahimiko wrote:
This question relates to absolute value. |6-4| =2 |3-8|= -5 ur going to turn that negative 5 into a postive 5 then 2-5= -3 so the answer will be -3
positive*

BettyIsAwesome:

TY TY TY

TogaHimiko:

np lol<3

BettyIsAwesome:

wait but why do u turn the negative 5 into a postive 5

TogaHimiko:

You turn the negative 5 into a positive 5 because the absolute value is regarding the actual number without having to do with the sign it has...An absolute value is an actual number itself... It can't be negative.. So lets say we have a question like what is the absolute value of |-4| it would be 4... 4 is the real number of -4..i think

BettyIsAwesome:

@togahimiko wrote:
You turn the negative 5 into a positive 5 because the absolute value is regarding the actual number without having to do with the sign it has...An absolute value is an actual number itself... It can't be negative.. So lets say we have a question like what is the absolute value of |-4| it would be 4... 4 is the real number of -4..i think
So your saying that absolute value is a non negative number?

TogaHimiko:

yes

TogaHimiko:

it is absolute Value is a non negative number.

BettyIsAwesome:

@togahimiko wrote:
yes
ty

TogaHimiko:

anytime<3